OpenStudy (anonymous):

antiderivitie of (x^2+1)^.5 with integral from 0 to 2 using fundamental theorms and if not possible, then using riemann sums to approximate

OpenStudy (amoodarya):

substitute x=sinh u so sinh^2u +1=cosh^2u

OpenStudy (anonymous):

what is sinh u?

OpenStudy (amoodarya):

ok if you don't know sinh use this substitute x=tan u so tan^2u+1=sec^2u

OpenStudy (anonymous):

we never did this? Are you sure that is the main approach?

OpenStudy (anonymous):

we learned fundamental theorm part 1 and 2 but never dealt with substitution. Maybe if it is not possible, should we estimate the integral using Riemann sums?

OpenStudy (anonymous):

The book's answer is 2.96